OpenStudy (anonymous):

The function rule C = 10n + 26 relates the number of people n who attend a small concert to the cost in dollars C of the concert. Make a table of the input/output pairs to show the cost if 27, 39, and 43 people attend. (3 points) (explain @boricua5

OpenStudy (anonymous):

@Seran

OpenStudy (anonymous):

cost 27 => =10*27+26 = 296 cost 39 => =10*39+26 = 416 cost 43 => =10*43+26 = 456

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

18. Explain how to write a function rule from the table below. Then write a function rule. x 2 4 6 y 1 0 –1 (3 points)

OpenStudy (anonymous):

@Seran

OpenStudy (anonymous):

let y = mx+c substitute x=2,y=1 1 = 2m+c sustitute x=4,y=0 0 = 4m+c solve these 2 eqation m=-1/2,c=2 so function is y=-0.5x+2

OpenStudy (anonymous):

19. A prototype rocket burns 159 gallons of liquid hydrogen fuel each second during a regular launch. Make a table that shows how many gallons of fuel are burned after 2, 3, 5, and 10 seconds. (3 points)

OpenStudy (anonymous):

@Vegiti